A normal population has a mean of 60 and a standard deviation of 3. You select a sample of 36.

Compute the probability the sample mean is (Round z values to 2 decimal places and final answers to 4 decimal places):
a. Less than 59.

Probability ___?


b. Between 59 and 61.

Probability _____?


c. Between 61 and 62.

Probability ____?


d. Greater than 62.

Probability ____?


Answer by Boreal(15235)

The standard error is 3/sqrt(36)=3/6=0.5
Less than 59 is z=(59-60)/0.5=-2
probability z < -2 is 0.0228
====
between 59 and 61 is the probability -2 ==========
between 61 and 62 is the probability z is between 2 and 4, and that is 0.0228.
=========
probability > 62 is z > (62-60)/0.5 or z >4, which is 0.0000317